OpenStudy (anonymous):

what is a front end estimation for 9.327+5.27+4.132

OpenStudy (gorv):

welcome to OS :)

OpenStudy (gorv):

first add front end 9+5+4=18

OpenStudy (gorv):

now round of decimal to nearest

OpenStudy (gorv):

0.327=0.3 0.27=0.3 0.132=0.1 =0.3+0.3+0.1=0.7 so front end = 18+0.7=18.7

OpenStudy (gorv):

let me give u any website if i find it

OpenStudy (gorv):

hmm... try this one http://www.mathgoodies.com/lessons/decimals/estimate_sums.html